### v3

@@ 3.2 Evaluation of the goodness of the model
Figure 3: The spatial binning scheme we used for comparing the velocity distributions of the model and data. We divide the model into $7\times 5$ spatial bins along $r\times\theta$, and calculate the likelihood of stars located within each bin by comparing to the velocity distributions of the model in the corresponding bin.
Figure 4: The model constraints on the parameters of the underlying DM distribution $\rho_{0}$, $r_{s}$, $\gamma$, $q_{\rm DM}$, and $p_{\rm DM}$. The three parameters determining the radial distribution, $\rho_{0}$, $r_{s}$, $\gamma$, are degenerate due to the lack of data points at $r<8$ kpc and $r>50$ kpc. However the 3D DM distribution within our data coverage ($r\lesssim 50$ kpc) are well constrained by our model, with a significant role played by the density distribution $\chi^{2}_{\rm den}$.
Figure 5: The enclosed mass profiles and rotation curve we obtained for the Milky Way. The solid line is the average of models within the $1\sigma$ confidence level, while the dashed curves are the upper and lower boundary.
We detail particular components of the DF: the spatial density distribution and velocity distributions at different positions, and we evaluate these for the model against the data.
We calculate the stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane, dividing it into $25\times 25$ bins within 50 kpc. The density in each bin is calculated by $N/(2\pi RdRdz)$ in units of $N/{\rm kpc}^{3}$. We calculate the density directly for both the observed data $\rho_{\rm data}$ and the model $\rho_{\rm model}$. The uncertainty of the data $d\rho_{\rm data}$ is assessed by bootstrapping the distances of stars along with their uncertainties. We calculate $\chi^{2}_{\rm den}$ by directly comparing the data and the model:
@@
To analyse velocity distributions, we categorise model particles into $N_{\rm bin}=7\times 5$ bins along the direction of $r$ and $\theta$ as shown in Fig. 3. We construct the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin $j$ and represent the distribution of each velocity component by a histogram using intervals of $M$, labelled ($v^{k}_{m,j}$, $h^{k}_{m,j}$), where $k$ corresponds to the velocities $v_{r}$, $v_{\phi}$, and $v_{\theta}$.
In bin $j$, we have $N_{j}$ stars, and for a star $i$ with a velocity and measurement uncertainty ($v^{k}_{i}$, $\sigma^{k}_{i}$) within that bin, we determine its likelihood compared to the model for each of the three velocity components as follows:
$P^{k}_{i,j}=\frac{1}{\sqrt{2\pi}\times\sigma^{k}_{i}}\times\frac{\sum_{m=1}^{M}h^{k}_{m,j}\times\exp[-(v^{k}_{i}-v^{k}_{m,j})^{2}/(2(\sigma^{k}_{i})^{2})]}{\sum_{m=1}^{M}(h^{k}_{m,j})}$ (14)
@@
Our data for integrating stellar orbits primarily cover the range $4<r<50$ kpc, meaning that we lack data for orbits starting at $r<4$ kpc. These orbits could significantly influence the DF of stars within $r<8$ kpc. Consequently, the estimated DF from the model remains incomplete for $r<8$ kpc, whereas it is more than $90\%$ complete for the range $8<r<50$ kpc. Thus, we exclude the five bins in $r=[4,8]$ kpc from our likelihood calculations. Although our observations for the southern hemisphere are incomplete, the 3082 stars available should conform to the same DF. We incorporate these stars into the northern spatial bins and include them in our likelihood calculations.
The likelihood across the three velocity components is merged to derive the constraints from the velocity distributions:
@@
Finally, we combine the $\chi^{2}$ values from both the density distribution and the velocity distributions to obtain
@@ 4.1 The 3D dark matter distribution of Milky Way
Figure 6: Comparison of the observational data and the empirical orbit superposed model: the stellar density distribution. The top panels from left to right are the 2D density distribution $\rho_{\rm star}$ in $R_{\rm gc}$ versus $z_{\rm gc}$ constructed by data, our best-fitting model ($q_{\rm DM}=0.92$, $p_{\rm DM}=0.80$), a model with a flatter DM halo, and a model with a rounder DM halo. The bottom panels are the relative error of the data, and the corresponding model residuals. The inset panel in the right shows the flattening of the stellar halo $q_{\rm star}\equiv z_{\rm gc}/R_{\rm gc}$ as a function of radius $R_{\rm gc}$ extracted from the iso-density contours of the 2D density distribution. The star symbols with errorbar are data, the solid red, light blue and black curves indicates the three models shown in the left. The red dashed curve indicate a model allowing $q_{\rm DM}$ and $p_{\rm DM}$ vary as a function of radius. The best-fitting model constructed stellar density distribution that matches the observational data well, while models with flatter or rounder DM halos constructed stellar halos either too flat or too round. The density distribution of the halo stars, especially the flattening $q_{\rm star}$ has strong constraints on the shape of the underlying DM halo.
@@
Figure 7: Comparison of the observational data and the empirical orbit superposed model: the 3D velocity distributions $v_{r}$, $v_{\phi}$, and $v_{\theta}$ in different spatial bins. We have divided the system into $7\times 5$ bins in $r$ versus $\theta$, but eliminating first row with $r<8$ kpc and only showing the odd number of columns here. In each sub-panel, the gray solid and dashed curves are the mean and $1\sigma$ uncertainty of the distribution constructed from observational data. The red, dark blue and light blue curves are constructed by models with different radial distribution of the underlying DM mass. The model with DM mass of $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=70$ kpc, $\gamma=1.0$ (red) matches the data well. The velocity distribution of stars, especially $v_{r}$ strongly constrains the radial distribution of the underlying DM distributions.
@@
We investigated the parameter space of $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ within the gravitational potential. A parameter grid is established with intervals of 0.1, 10 kpc, 0.02, 0.04, and 0.02 for the five parameters, respectively. An iterative approach is used to find the optimal models. The process begins with an initial model, and after developing initial models, iterative refinement follows. During each iteration, we identify the optimal models using the criterion $\chi^{2}-\min(\chi^{2})<\chi^{2}_{s}$, where $\chi^{2}_{s}=100$. We then generate new models by walking two steps in every direction of the parameter grid of each optimal model. This approach guides the search towards the lower $\chi^{2}$ within the parameter grid and halts when the model with the minimum $\chi^{2}$ is identified. The relatively high value of $\chi^{2}_{s}$ is selected to ensure that all models within the confidence level of $1\sigma$ are considered before the end of the iteration. Ultimately, we determine the best-fitting models by achieving the minimum $\chi^{2}_{\rm tot}$.
@@
The $1\sigma$ confidence level is determined by bootstrapping. We take the gravitational potential of the best-fitting model and perturb the position and velocity of observed stars with their uncertainty for 100 trials. In each trial, we recompute the model and evaluate the resulting $\chi^{2}_{\rm den}$, $\chi^{2}_{\rm vdis}$ and $\chi^{2}_{\rm tot}$. The standard deviation of these 100 values of $\chi^{2}$ is taken as the $1\sigma$ confidence level of the model $\Delta\chi^{2}_{1\sigma}$.
@@
The parameter grid we explored for the Milky Way is illustrated in Fig.4 with about 15,000 models calculated. Both the density distribution $\chi^{2}_{\rm den}$ and the velocity distributions $\chi^{2}_{\rm kin}$ yield consistent results of the gravitational potential parameters, with $\chi^{2}_{\rm den}$ playing a key role in determining the DM shape parameters $p_{\rm DM}$ and $q_{\rm DM}$. We establish the DM halo shape parameters as $q_{\rm DM}=0.92\pm 0.08$ and $p_{\rm DM}=0.8\pm 0.2$, a nearly oblate DM halo, but with the long-intermediate axis plane vertically aligned with the stellar disk ($q_{\rm DM}>p_{\rm DM}$) that is strongly preferred by our model.
@@
The three parameters determining the radial distribution, $\rho_{0}$, $r_{s}$, $\gamma$, are still degenerate. However, the enclosed DM mass profile within the data coverage is reasonably constrained. We determine the DM mass to be $M_{\rm DM}(<50{\rm kpc})=(5.4\pm 1.6)\times 10^{11}\,M_{\odot}$. The circular velocity at the solar location ($r=8.2$ kpc (Bland-Hawthorn2016)) is found to be $V_{c}=220_{-23}^{+29}$ km/s, which agrees well with previous findings (McMillan2017MNRAS.465...76M). Our derived enclosed mass profile and rotation curve for the Milky Way are presented in Fig.5.
@@ 4.2 The best-fitting models
To illustrate how well the models match the data, we show the stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane in Fig.6. The stellar density distributions of the data and model are constructed in the same manner. We further extract the flattening of the halo stars $q_{\rm star}$ along the iso-density contours of $\log(\rho_{\rm star})=[-3,-2,-1,0,1,2,3]$ and compare that from the data and the model. As shown in Fig.6, the model with an underlying DM halo of $q_{DM}=0.92$, $p_{\rm DM}=0.80$ constructed a stellar halo density that matches the data well. Variations in DM halo shape yield significant deviations in the stellar density distribution of the orbit superposition model from the actual data, especially the flattening $q_{\rm star}$ at different radii. In other words, the density distribution of the halo stars, especially the flattening $q_{\rm star}$ has strong constraints on the shape of the underlying DM halo; our results on the 3D shape of DM halo are highly data-driven.
To analyse velocity distributions, we categorised model particles into bins $N_{\rm bin}=7\times 5$ in $r$ and $\theta$. The divisions are $r=[4,8,10,12,15,20,30,50]$ kpc and $\theta=[0,15,30,45,60,90^{\circ}]$. We then construct the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin for the data and model, as illustrated in Fig.7. Likewise, the model with DM mass of $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=70$ kpc, $\gamma=1.0$ (red) matches the data well, while an altered density profile results in notable differences in the model’s velocity distributions. The velocity distribution of stars, especially $v_{r}$ strongly constrains the radial distribution of the underlying DM distributions.
Also note that our best-fitting model, which constructed both stellar density and the $3D$ velocity distributions, agrees with the data generally well, supporting our hypothesis of a ‘stationary system’.
@@ 4.3 Models with variable $p_{\rm DM}(r)$, $q_{\rm DM}(r)$
The best-fitting models with constant $p_{\rm DM}$, $q_{\rm DM}$ generally succeed in replicating the full DF of the data. There is a small deviation between the 2D stellar density distribution constructed from the data and the model. This deviation is not statistically significant, but it becomes noticeable when we evaluate the flattening of the stellar halo ($q_{\rm star}$) along the iso-density contours, as shown in the right panel of Fig. 6.
We further create a set of models for the Milky Way with $p_{\rm DM}(r)$, $q_{\rm DM}(r)$ following equations 9 to 12. We fix $r_{q}=20$ kpc and allow $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$ and $q_{\rm out}$ to be free parameters. The parameters determining the radial DM profile $\rho_{0}$, $r_{s}$, and $\gamma$ do not degenerate significantly with the 3D shape. To optimise the efficiency of the model, we fix them at $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.3$, $r_{s}$=40 kpc, and $\gamma=1.4$. We created about 2000 models to explore the parameter space of $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$ and $q_{\rm out}$. We obtained models with $q_{\rm in}<q_{\rm out}$ and $p_{\rm in}>p_{\rm out}$ preferred, although statistically not significant. The flattening of the stellar halo ($q_{\rm star}$) of a best-fitting model is shown in Fig.6, which indeed matches the data slightly better. We extract $p_{\rm DM}$ and $q_{\rm DM}$ at $r=15,30,50$ kpc from each model and will compare these with cosmological simulations in the next section.

### v4

@@ 3.2 Evaluation of the goodness of the model
Figure 3: Spatial binning scheme we used for comparing the velocity distributions of the model and data. We divided the model into $7\times 5$ spatial bins along $r\times\theta$, and calculated the likelihood of stars located within each bin by comparing to the velocity distributions of the model in the corresponding bin.
Figure 4: Model constraints on the parameters of the underlying DM distribution, $\rho_{0}$, $r_{s}$, $\gamma$, $q_{\rm DM}$, and $p_{\rm DM}$. The three parameters determining the radial distribution, $\rho_{0}$, $r_{s}$, and $\gamma$, are degenerate due to the lack of data points at $r<8$ kpc and $r>50$ kpc. However the 3D DM distribution within our data coverage ($r\lesssim 50$ kpc) are well constrained by our model, with a significant role played by the density distribution, $\chi^{2}_{\rm den}$.
Figure 5: Enclosed mass profiles and rotation curve that we obtained for the Milky Way. The solid line is the average of models within the $1\sigma$ confidence level, while the dashed curves are the upper and lower boundary.
We detail particular components of the DF: the spatial density distribution and velocity distributions at different positions, and we evaluate these for the model against the data. We calculated the stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane, dividing it into $25\times 25$ bins within 50 kpc. The density in each bin was calculated by $N/(2\pi RdRdz)$ in units of $N/{\rm kpc}^{3}$. We calculated the density directly for both the observed data, $\rho_{\rm data}$, and the model, $\rho_{\rm model}$. The uncertainty of the data, $d\rho_{\rm data}$, was assessed by bootstrapping the distances of stars along with their uncertainties. We calculated $\chi^{2}_{\rm den}$ by directly comparing the data and the model:
@@
To analyse velocity distributions, we categorised model particles into $N_{\rm bin}=7\times 5$ bins in the direction of $r$ and $\theta$ as is shown in Fig. 3. We constructed the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin, $j$, and represent the distribution of each velocity component by a histogram using intervals of $M$, labelled ($v^{k}_{m,j}$, $h^{k}_{m,j}$), where $k$ corresponds to the velocities $v_{r}$, $v_{\phi}$, and $v_{\theta}$.
In bin $j$, we have $N_{j}$ stars, and for a star, $i$, with a velocity and measurement uncertainty ($v^{k}_{i}$, $\sigma^{k}_{i}$) within that bin, we determined its likelihood compared to the model for each of the three velocity components as follows:
$P^{k}_{i,j}=\frac{1}{\sqrt{2\pi}\times\sigma^{k}_{i}}\times\frac{\sum_{m=1}^{M}h^{k}_{m,j}\times\exp[-(v^{k}_{i}-v^{k}_{m,j})^{2}/(2(\sigma^{k}_{i})^{2})]}{\sum_{m=1}^{M}(h^{k}_{m,j})}.$ (14)
@@
Our data for integrating stellar orbits primarily cover the range $4<r<50$ kpc, meaning that we lack data for orbits starting at $r<4$ kpc. These orbits could significantly influence the DF of stars within $r<8$ kpc. Consequently, the estimated DF from the model remains incomplete for $r<8$ kpc, whereas it is more than $90\%$ complete for the range $8<r<50$ kpc. Thus, we excluded the five bins in $r=[4,8]$ kpc from our likelihood calculations. Although our observations for the southern hemisphere are incomplete, the 3082 stars available should conform to the same DF. We incorporated these stars into the northern spatial bins and included them in our likelihood calculations.
The likelihood across the three velocity components was merged to derive the constraints from the velocity distributions:
@@
Finally, we combined the $\chi^{2}$ values from both the density distribution and the velocity distributions to obtain
@@ 4.1 The 3D dark matter distribution of Milky Way
Figure 6: Comparison of the observational data and the empirical orbit superposed model: the stellar density distribution. The top panels from left to right are the 2D density distribution $\rho_{\rm star}$ in $R_{\rm gc}$ versus $z_{\rm gc}$ constructed by data, our best-fitting model ($q_{\rm DM}=0.92$, $p_{\rm DM}=0.80$), a model with a flatter DM halo, and a model with a rounder DM halo. The bottom panels show the relative error of the data, and the corresponding model residuals. The inset panel in the right shows the flattening of the stellar halo, $q_{\rm star}\equiv z_{\rm gc}/R_{\rm gc}$, as a function of the radius, $R_{\rm gc}$, extracted from the iso-density contours of the 2D density distribution. The star symbols with the error bar are data. The solid red, light blue, and black curves indicate the three models shown on the left. The dashed red curve indicates a model allowing $q_{\rm DM}$ and $p_{\rm DM}$ to vary as a function of radius. The best-fitting model constructed stellar density distribution that matches the observational data well, while models with flatter or rounder DM halos constructed stellar halos that are either too flat or too round. The density distribution of the halo stars, especially the flattening $q_{\rm star}$, has strong constraints on the shape of the underlying DM halo.
@@
Figure 7: Comparison of the observational data and the empirical orbit superposed model: the 3D velocity distributions $v_{r}$, $v_{\phi}$, and $v_{\theta}$ in different spatial bins. We divided the system into $7\times 5$ bins in $r$ versus $\theta$, but eliminating the first row with $r<8$ kpc and only showing the odd number of columns here. In each sub-panel, the solid and dashed grey curves are the mean and $1\sigma$ uncertainty of the distribution constructed from observational data. The red, dark blue, and light blue curves were constructed by models with different radial distributions of the underlying DM mass. The model with a DM mass of $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=70$ kpc, and $\gamma=1.0$ (red) matches the data well. The velocity distribution of stars, especially $v_{r}$, strongly constrains the radial distribution of the underlying DM distributions.
@@
We investigated the parameter space of $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ within the gravitational potential. A parameter grid was established with intervals of 0.1, 10 kpc, 0.02, 0.04, and 0.02 for the five parameters, respectively. An iterative approach was used to find the optimal models. The process began with an initial model, and after developing initial models, iterative refinement followed. During each iteration, we identified the optimal models using the criterion $\chi^{2}-\min(\chi^{2})<\chi^{2}_{s}$, where $\chi^{2}_{s}=100$. We then generated new models by walking two steps in every direction of the parameter grid of each optimal model. This approach guided the search towards the lower $\chi^{2}$ within the parameter grid and halted when the model with the minimum $\chi^{2}$ was identified. The relatively high value of $\chi^{2}_{s}$ was selected to ensure that all models within the confidence level of $1\sigma$ were considered before the end of the iteration. Ultimately, we determined the best-fitting models by achieving the minimum $\chi^{2}_{\rm tot}$.
@@
The $1\sigma$ confidence level was determined by bootstrapping. We took the gravitational potential of the best-fitting model and perturbed the position and velocity of observed stars with their uncertainty for 100 trials. In each trial, we recomputed the model and evaluated the resulting $\chi^{2}_{\rm den}$, $\chi^{2}_{\rm vdis}$, and $\chi^{2}_{\rm tot}$. The standard deviation of these 100 values of $\chi^{2}$ was taken as the $1\sigma$ confidence level of the model $\Delta\chi^{2}_{1\sigma}$.
@@
The parameter grid we explored for the Milky Way is illustrated in Fig.4 with about 15,000 models calculated. Both the density distribution, $\chi^{2}_{\rm den}$, and the velocity distributions, $\chi^{2}_{\rm kin}$, yield consistent results of the gravitational potential parameters, with $\chi^{2}_{\rm den}$ playing a key role in determining the DM shape parameters, $p_{\rm DM}$ and $q_{\rm DM}$. We established the DM halo shape parameters as $q_{\rm DM}=0.92\pm 0.08$ and $p_{\rm DM}=0.8\pm 0.2$, a nearly oblate DM halo, but with the long-intermediate axis plane vertically aligned with the stellar disc ($q_{\rm DM}>p_{\rm DM}$) that is strongly preferred by our model.
@@
The three parameters determining the radial distribution, $\rho_{0}$, $r_{s}$, and $\gamma$, are still degenerate. However, the enclosed DM mass profile within the data coverage is reasonably constrained. We determined the DM mass to be $M_{\rm DM}(<50{\rm kpc})=(5.4\pm 1.6)\times 10^{11}\,M_{\odot}$. The circular velocity at the solar location ($r=8.2$ kpc (Bland-Hawthorn and Gerhard, 2016)) was found to be $V_{c}=220_{-23}^{+29}$ km/s, which agrees well with previous findings (McMillan, 2017). Our derived enclosed mass profile and rotation curve for the Milky Way are presented in Fig.5.
@@ 4.2 The best-fitting models
To illustrate how well the models match the data, we show the stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane in Fig.6. The stellar density distributions of the data and model were constructed in the same manner. We further extracted the flattening of the halo stars, $q_{\rm star}$, along the iso-density contours of $\log(\rho_{\rm star})=[-3,-2,-1,0,1,2,3]$ and compared that from the data and the model. As is shown in Fig.6, the model with an underlying DM halo of $q_{DM}=0.92$, $p_{\rm DM}=0.80$ constructed a stellar halo density that matches the data well. Variations in the DM halo shape yield significant deviations in the stellar density distribution of the orbit superposition model from the actual data, especially the flattening $q_{\rm star}$ at different radii. In other words, the density distribution of the halo stars, especially the flattening $q_{\rm star}$, places strong constraints on the shape of the underlying DM halo; our results for the 3D shape of DM halo are highly data-driven.
To analyse velocity distributions, we categorised model particles into bins of $N_{\rm bin}=7\times 5$ in $r$ and $\theta$. The divisions are $r=[4,8,10,12,15,20,30,50]$ kpc and $\theta=[0,15,30,45,60,90^{\circ}]$. We then constructed the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin for the data and model, as is illustrated in Fig.7. Likewise, the model with DM mass of $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=70$ kpc, $\gamma=1.0$ (red) matches the data well, while an altered density profile results in notable differences in the model’s velocity distributions. The velocity distribution of stars, especially $v_{r}$, strongly constrains the radial distribution of the underlying DM distributions. We also note that our best-fitting model, which constructed both stellar density and the $3D$ velocity distributions, agrees with the data generally well, supporting our hypothesis of a ‘stationary system’.
@@ 4.3 Models with variable $p_{\rm DM}(r)$, $q_{\rm DM}(r)$
The best-fitting models with constant $p_{\rm DM}$, $q_{\rm DM}$ generally succeed in replicating the full DF of the data. There is a small deviation between the 2D stellar density distribution constructed from the data and the model. This deviation is not statistically significant, but it becomes noticeable when we evaluate the flattening of the stellar halo ($q_{\rm star}$) along the iso-density contours, as is shown in the right panel of Fig. 6.
We further created a set of models for the Milky Way with $p_{\rm DM}(r)$, $q_{\rm DM}(r)$ following equations 9 to 12. We fixed $r_{q}=20$ kpc and allowed $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$, and $q_{\rm out}$ to be free parameters. The parameters determining the radial DM profile, $\rho_{0}$, $r_{s}$, and $\gamma$, do not degenerate significantly with the 3D shape. To optimise the efficiency of the model, we fixed them at $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.3$, $r_{s}$=40 kpc, and $\gamma=1.4$. We created about 2000 models to explore the parameter space of $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$, and $q_{\rm out}$. We obtained models in which $q_{\rm in}<q_{\rm out}$ and $p_{\rm in}>p_{\rm out}$ are preferred, although statistically not significant. The flattening of the stellar halo ($q_{\rm star}$) of a best-fitting model is shown in Fig.6, which indeed matches the data slightly better. We extracted $p_{\rm DM}$ and $q_{\rm DM}$ at $r=15,30,50$ kpc from each model and compare these with cosmological simulations in the next section.